OpenStudy (anonymous):

Can someone please help? What is a basis for the column space of matrix A = [ 1,2,0,1; 0,1,1,0; 1,2,0,1]? Why?

OpenStudy (phi):

use elimination to find the pivot columns.

OpenStudy (anonymous):

I did and got that x1 and x3 are pivot columns, but according to the solution they are x1 and x3, and don't know why...

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 2 & 0 & 1 \\ 0 & 1 & 1 & 0\\ 1&2&0&1\end{matrix}\right]\]

OpenStudy (anonymous):

R1-R3

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 2 & 0 & 1 \\ 0 & 1 & 1 & 0\\ 0&0&0&0\end{matrix}\right]\]

OpenStudy (anonymous):

that's right... that's what I have... so I said the first two columns were the basis, but supposedly it is the first and third...

OpenStudy (phi):

You are not wrong. However, you could use any 2 independent columns.

OpenStudy (phi):

clear col 4 equals col 1, so you can not use col1 and col4

OpenStudy (anonymous):

thanks

OpenStudy (phi):

btw, dumb book!

OpenStudy (anonymous):

agree!